OpenStudy (anonymous):

there are 12 nickels, dimes, and quarters whose total value is $1.55. if the number of nickels exceeds the numbers of quarters by 1, how many coins of each kind are there?

OpenStudy (anonymous):

.5x+.10y+.25z=1.55 or 5x+10y+25z=155

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Let n, d and q equal the number of nickles, dimes and quarters respectively. Solve the following for n d and q.\[\{5n+10d+25q=155,n+d+q=12,n=q+1\}\]\[\{n=5,d=3,q=4\} \]

OpenStudy (anonymous):

Original posting had an error in the third equation.

OpenStudy (anonymous):

i understand now... the book has examples but its messy, thank you

OpenStudy (anonymous):

how did you get 12?

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

In this class of problems I found it better to deal with pennies. Thank you for the medal.

OpenStudy (anonymous):

easier